Tony was arranging the CDs for the yard sale. So far he had five different groups. He had put 1/8 of the CDs in the jazz section, 1/4 into the rap section, 3/8in the rock section and 3/16 in the pop section. He put 3 CDs in the classical section. How many CDs did Tony arrange, and how many were in each section.
Answer by mananth(16946)

Tony was arranging the CDs for the yard sale.
He had put 1/8 of the CDs in the jazz section x/8
1/4 into the rap section x/4
,3/8in the rock section3x/8
and 3/16 in the pop section.3x/16
He put 3 CDs in the classical section. 3
x/8 +x/+3x/8+3x/16 +3 = x
multiply each fraction by 16
2x+4x+6x+3x+48 = 16x
x=48
x/8 =6 Jazz
x/4= 12 rap
3x/8 = 18 Rock
3x/16 =9 Pop
+3 classical